Labor Economics

9. Phils wage increases and Phil responds by working more hours. Which of the following could explain this? a. Phils substitution effect is larger than his income effect b. Phil views leisure as an inferior good. c. Phils income effect is larger than his substitution effect d. Both b and c f. Phil died, but was brought back to life with an alien serum in a plot twist no one saw because they gave up after episode 5. e. Both a and b 10. Where do labor supply curves come from? a. Our professor just makes them up b. Individuals making tradeoffs between leisure and income c. The value of the marginal product of labor d. The inverse of a workers highest attainable indifference curve e. When a mommy labor supply curve and a daddy labor supply curve love each other very much... 11. Given the following data, what is the labor force participation rate? Populatiorn Civilian, non-institutionalized population, age 16 and over Employed persons Unemployed persons 200 million 150 million 100 million 50 million a. 66.7% b. 75% с. 80% d. 84% e. 100% 12. In problem 11, how many NILFs were there? a. none e. I dont know; I was expecting some kind of joke about your mom b. 50 million c. 100 milliond. 200 million 13. If discouraged workers all decided to start looking for work, what would happen to the labor force participation rate? (Assume none of them are actually going to find iobs.) a. It would increase b. It would decrease c. No change. d. Wed need more information14. Marias employer is considering her for a training program that will cost $3 per hour. Her current VMPL is $15 per hour and will rise to $30 upon completion of the program. The training would make her more productive for her employer, but also for any other employer. Of the following, Marias training and post-training wage, respectively, will most likely be a. $15, S30 b. S12; $15 c. S15; $16 d. $12; S30 e. $12, $27 15. T/F The total cost of a college education is the sum of all the money you spend while in college.
Answer #1

Solution:

9. We will first see how the substitution and income effects take place under such scenario.

With increase in Phil's wage, the cost of leisure increases. Since, leisure becomes expensive, it's demand will decrease, thereby increasing the labor supply. This is the substitution effect. Also, with increase in the wage, labor income increases (for initial hours of labor supplied). With assumption that leisure is a normal good, demand for leisure increases (so labor supply falls). This is the income effect.

Since, with increase in Phil's wage, as we are given Phil increases his labor supply (that is, now work for more hours), clearly, from above reasoning we can say that substitution effect must have dominated the income effect. Also, in our reasoning above, we have already taken/considered leisure as a normal good, and not inferior good. So, the correct option is a) Phil's substitution effect is larger than his income effect.

10. We answer this question by method of elimination (of wrong options). Options (a) and (e) are undoubtedly futile options, so eliminating them. Option (c) talks about the marginal product of labor, which helps in determining the productivity of labor for a firm, and thus, mark the demand side of labor market and not the supply side. Option (d) is a wrong one as a higher indifference curve will provide worker with more utility, and the worker will always try to attain that, hence it's inverse will give falsified result.

Correct option is (b) Individuals making tradeoffs between leisure and income, which is same as saying tradeoffs between leisure and labor (as providing labor will generate income). With changing wages, a person has to always decide how to distribute the limited time endowment between the two: labor and leisure, which in turn helps us generate the labor supply curve.

11. We are given that total population (P) is 200 million, of which non-institutionalized population (NIP) (people who are allowed to work by law) = 150 million. Among these non-institutionalized people, 100 million are employed (E) persons and 50 million are unemployed (UE) persons.

Labor force (LF) include people who are employed and unemployed.

So, accordingly, LF = E + UE = 100 + 50 = 150 million

Labor force participation rate = (LF/NIP)*100 = (150/150)*100 = 100

So, LFPR = 100%, making option (e) 100% as the correct option.

12. NILF stands for Not-In-Labor Force. The civilian or non-institutionalized persons can be categorized in 3: Employed persons, Unemployed persons and NILFs. The first two together make up the labor force (as seen in question 11), and the last ones include persons who are able (by law, as they have crossed the required age of 16), but not actively looking for a job.

So, NILF = NIP - LF

With NIP = LF (as seen in question 11), NILF = 0.

So, there were 0 (or no) NILFs, making option (a) none, the correct one.

13. Discouraged workers are the ones who do not actively look for jobs (so they are part of NILF). When they actively start looking for the job, they become a part of labor force. (Since, they actually do not find the jobs, they'll precisely be part of unemployed population, and not the employed one, but in any case, they'll enter the labor force). So, with such transition, even though number of non-institutionalized persons remain unchanged, LF increases.

LFPR = LF/NIP. With increased LF (numerator value) and same NIP (denominator value), the LFPR would increase. So, correct option is (a) It would increase.

14. In labor market, wage equals the VMPL (Value of marginal product of labor). So, Maria's wage during training period = $15 per hour. But the training will cost $3 per hour as well, so ultimately during this period, wage = 15 - 3 = $12.

Post-training, as mentioned, VMPL rises to $30, so post-training wage = $30. Note that Maria was enrolled into a training program initially, but now in post-training, she has no such costs to incur and thus can retain her entire wage. Her final wage, thus, still is $30.

So, correct option is (d) $12; $30

15. FALSE. The most basic example that can be stated here to contradict the given statement is the cost you incur in entertainment and fun. So, if you buy passes to a show in your college, or eat too often say in the college canteen, etc., this is part of money you spend while in college, but this is not cost under college education. College education cost mainly includes the tuition fees, the hostel charge (if you're staying in one) and such unavoidable costs. We can therefore claim that given statement is false.